OpenStudy (anonymous):

is it possible to separate this differential equation dy/dx - ycotx = x or could the question be wrong?

OpenStudy (amistre64):

you might have to employ a different tactic

OpenStudy (amistre64):

get it in the form of y/x i believe

OpenStudy (amistre64):

then replace it with a "v" or "z"

OpenStudy (amistre64):

its not a technique i do enough to recall the details; but i think thats the brunt of it

OpenStudy (amistre64):

\[y' = x+ y\ cot(x)\] \[y' = 1+ \frac{y}{x}\ cot(x)\] z = y/x z' = (xy' - x'y) x^2 and solve for y' and sub it in

OpenStudy (amistre64):

that make any sense?

OpenStudy (amistre64):

\[y' = 1+ \frac{y}{x}\ cot(x)\] \[z'x +z = 1+ z\ cot(x)\] \[z'x = 1+ z\ cot(x)-z\] \[z' = \frac{1+ z\ cot(x)-z}{x}\] if i remember it right

OpenStudy (anonymous):

Yes it makes sense but the question says to solve the differential equation and i'm not sure how introducing z helps?

OpenStudy (amistre64):

z is just a substitution for "y/x" to make it easier to seperate and solve

OpenStudy (amistre64):

or its spose to :)

OpenStudy (amistre64):

i almost had it right :) since z=y/x; then y = zx, and derive for y'

OpenStudy (amistre64):

since z is lousy in the font ill use v \[v=\frac{y}{x}\] \[y=vx\] \[y'=v+x\frac{dv}{dx}'\]

OpenStudy (amistre64):

heres a good explanation if we can use it: http://www.youtube.com/watch?v=UpLQUGBznE4&feature=related

OpenStudy (anonymous):

Thanks!

OpenStudy (amistre64):

youre welcome, hope it helps ..